Coupon bond that pays interest of 4.69% annually has a par value of $1,000 matures in 20 years, and is selling today at $850. Actual yield?

The company than downgrades from investment grade to speculative grade and the new required yield is equal to 8%. What's the bond price now?
Answer #1

Coupon =4.69%*1000 =46.90
Number of years =20
Par Value =1000
Yield using excel function =RATE(20,46.90,-850,1000) =6%

New required rate =8%
Price of bond =PV of coupons+PV of Par Value =46.90*((1-(1+8%)^-20)/8%+1000/(1+8%)^20 =675.02